OpenStudy (anonymous):

√(80)+√5:(108)-√(5) simplify and show steps please

OpenStudy (anonymous):

im struggling with this one - is that a ratio sign (:)?

OpenStudy (anonymous):

that nothing

OpenStudy (anonymous):

oh ok - thats what was confusing me - ill try again

OpenStudy (anonymous):

\[\sqrt[5]{108}\]

OpenStudy (anonymous):

= 4√5 + 108√5 - 5 = 112√5 - 5

OpenStudy (anonymous):

??

OpenStudy (anonymous):

how did u get 112√5 - 5

OpenStudy (anonymous):

5√108 not 108√5

OpenStudy (anonymous):

well √80 = √16 x √5 = 4√5 then i expanded the brackets √5 x 108 = 108√5 + √5 x √5 = 108√5 + 5

OpenStudy (anonymous):

5√108 = 108√5 - the order does not matter with multiplication

OpenStudy (anonymous):

3 x 4 = 4 x 3 : - both = 12

OpenStudy (anonymous):

oh i see - sorry

OpenStudy (anonymous):

the question says √5 (108 - √5) - what should it be?

OpenStudy (anonymous):

5√108 not 108√5

OpenStudy (anonymous):

√(80)+5√(108)-√(5)

OpenStudy (anonymous):

ok - thats different

OpenStudy (anonymous):

= 4√5 + 5*6√3 -√5 = 3√5 + 30√3 = 3(√5 + 10√3)

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

how did u get 3√5